OpenStudy (anonymous):

Can you solve these ? 2(3x+5)=40 6-X-X=18

OpenStudy (anonymous):

Distribute the 2.

OpenStudy (one098):

Distribute the 2, to the 3x and the 5.

OpenStudy (anonymous):

The first one is x=5, because... Simplifying 2(3x + 5) = 40 Reorder the terms: 2(5 + 3x) = 40 (5 * 2 + 3x * 2) = 40 (10 + 6x) = 40 Solving 10 + 6x = 40 Solving for variable 'x'. Move all terms containing x to the left, all other terms to the right. Add '-10' to each side of the equation. 10 + -10 + 6x = 40 + -10 Combine like terms: 10 + -10 = 0 0 + 6x = 40 + -10 6x = 40 + -10 Combine like terms: 40 + -10 = 30 6x = 30 Divide each side by '6'. x = 5 Simplifying x = 5

OpenStudy (anonymous):

~The best website ever

OpenStudy (one098):

Alright then!

OpenStudy (anonymous):

But that's if you're trying to find what x equals... >.> yerpaderp

OpenStudy (anonymous):

thank you soo much :)

OpenStudy (anonymous):

Yeah, no problem... c: But you should read what I posted, it helped me to do it this way so much when I needed help in this c:

OpenStudy (anonymous):

yeah i read it, you broke it down good. i understood it.